A car with a mass of 960 kg is lifted with a hydraulic press. The surface of piston 1 is 20cm² and the surface of piston 2 is 0.50m² . Piston 1 has a stroke length of 20cm.

a) What is the absolute pressure of the oil in the press?

b) What force is needed each time to push piston 1 down?

c) How many strokes are needed to lift the car 10cm?

Answer: a)18.8kPa b) 37.7N c) 125
Answer #1

Abs pressure = mg/A =960*9.8/0.5= 18.8Kpa

Force need to push = Pa *A = 18.8*0.002 kN =37.6N

Weight of car = 960*9.8= 9408 N

So since in force 37.6 N we are lifting 20 cm so to lift 9408 N 10 cm we need to press 9408/2*37.6 = 125 times